# Dynamics of systems with varying number of particles: from Liouville equations to general master equations for open systems

### v1

@@ II.2 The Bergman-Lebowitz equation of open systems
Bergmann and Lebowitz, in a seminal paper [34], have proposed an equation based on the extension of the Liouville equation that models open systems embedded in a reservoir of particles and energy. The essence of the model is to consider an impulsive, Markovian interaction between the reservoir and the system; the reservoir is considered stationary and not influenced by the changes occurring in the system, thus the thermodynamic state point of the reservoir is fixed. The interaction between the system and the reservoir is modeled as a discontinuous transition of the system from a state with particles () to one with particles (), where corresponds to the canonical variables (position and momenta). The change of the state of the system state is described by a time-independent Markovian transition kernel, . The kernel, , corresponds to the probability per unit time that the system at makes a transition to as a result of the interaction with the reservoir. The probability density, , in some point of the phase space is then regulated by an extension of the Liouville equation:
@@
The first line of eq. 16 corresponds precisely to the Liouville equation for the particle system, while the second line establishes a coupling with the reservoir(s). This last term is composed of negative or positive contributions corresponding either to the outflux/loss of probability from the current state or to the influx/gain of probability into the current state, respectively.
@@
it follows that the stationary Grand Ensemble is the Grand Canonical ensemble; here where is the Boltzmann constant and the temperature, while is the chemical potential of the system. The key difference between the approach of section II.1 and the approach of Bergman and Lebowitz lies in the assumption about the reservoir and the corresponding term of system-reservoir exchange in the equation. The model of section II.1 does not require any assumption about the reservoir, but directly integrates its particle’s degrees of freedom. As a consequence, the system-reservoir coupling term is explicitly written in terms of particle quantities without any stochastic assumption. In the model shown here, the reservoir is modeled a priori without any explicit link to its particle resolution and the system-reservoir coupling term is modeled, as a consequence, with a probabilistic process. The model of Bergman and Lebowitz has been very important for the numerical implementation of the open system approach AdResS in an intermediate step. Such model, differently from the model of section II.1 (as reported above in section II.1.1) could not provide an explicit numerical receipt for the system-reservoir coupling, nevertheless it allowed for a physical qualitative interpretation of the coupling conditions of the AdResS code [2, 16].
III Master equations for reaction-diffusion processes
@@
Molecular dynamics are limited to the study of complex biochemical phenomena at the scale of living cells. While the most sophisticated molecular simulations can achieve simulations of a few macromolecules in the scale of micro- or milliseconds when being optimistic, biochemical processes at these scales often involve thousands or millions of macromolecules and occur over timescales of seconds. Moreover, the chemical events at the molecular scale happen at a much faster scale than those relevant to life processes, and thus their detailed molecular kinetics do not play a key role in the dynamics. It is thus appropriate to consider particle-based reaction-diffusion models (fig. 3), where molecules are represented as particles undergoing random motion due to thermal fluctuations of the solvent (diffusion), and chemical reactions due to instantaneous reaction events, which often occur after chancy encounters between two or more molecules. The relevant features of the molecular scale are captured in the diffusion coefficients and the reaction rate functions.
@@
Particle-based reaction-diffusion models are often the standard to model biochemical processes at the scale of living cells. At these scales, the so-called chemical diffusion master equation (CDME) provides a probabilistic model in terms of the numbers of particles of each of the chemical species involved, as well as their spatial configuration [8, 11, 15]. One of the original objectives of developing the CDME was to serve as an underlying ground model upon which one can construct particle-based reaction-diffusion simulations consistently. Moreover, one can in principle recover most other reaction-diffusion models as limits of the CDME, e.g. in the well-mixed limit, one recovers the well-known chemical master equation (CME) [20, 23, 42, 45]; in the thermodynamic limit, one recovers reaction-diffusion PDEs; and when doing spatial discretizations, one recovers the reaction-diffusion master equation [28] or the spatiotemporal master equation [48]. Thus, it serves as a unifying framework from which one can develop numerical schemes that are consistent across scales.
@@
One of the main virtues of the CDME is its capability to handle systems with varying numbers of particles while maintaining spatial resolution. This is an inherent characteristic of the CDME since reaction events often change the number of particles in the system. In this section, we will overview the CDME for a simple example, and we will show how it extends when the diffusion term is replaced by Langevin dynamics.
a. b.
@@
III.1 Chemical diffusion master equation: coupling diffusion with reaction processes
@@
As a starting point, we follow [11] and consider a system with a varying number of particles, all corresponding to one chemical species, enclosed in a finite domain (reflecting boundaries in the boundaries of ). The configuration of the system is given by the number of particles and their positions, so its probability distribution is given as an ordered family of probability density functions:
@@
The generator of the CDME is decomposed into the reaction and diffusion components: is the diffusion operator and corresponds to the reaction operator for the th reactions. Each reaction operator can be split into two contributions
representing the gain or loss of probability respectively due to all possible reactions in a given configuration.
To exemplify the formulation of the CDME for a reaction involving only one chemical species, we focus on a reaction that models social friction in population dynamics (fig. 3b.)
@@
In the absence of drift and assuming isotropic diffusion, would simply be the Laplacian, with a scalar diffusion constant . Considering this structure and assuming there are no reactions, the CDME would correspond to an infinite family of uncoupled Fokker–Planck equations, where each member of the family corresponds to a different number of particles.
@@
where the notation omitted the time dependence of for simplicity. The extension to multiple species is reported in section III.1.1. The CDME for general one-species and multiple-species reactions were formulated in detail in [11, 8].
III.1.1 Multiple species extension
@@
with rate function , where and are the locations of one pair of reactants and is the location of the product. The stochastic dynamics of the system is described in terms of the distributions , where indicate the numbers of , , and particles, respectively, and indicates the positions of the particles, of the particles, and of the particles. The normalization condition eq. 19 generalizes to multiple species
@@
where to simplify notation, the dependence of on time has been omitted, as well as its dependence on the positions if clear from context. The first line corresponds to the diffusion of each chemical species, the second line to the gain term and the last one to the loss.
III.2 Langevin dynamics with varying number of particles
There is a close mathematical link between section II and section III.1; the resulting equations from both approaches describe the dynamics of densities in the phases space of systems with varying number of particles. In section III.1, we introduced the diffusion operator, as well as the densities, only dependent on the position of the particles. However, one can in principle define these operators on more general dynamics. For instance, the state of the system can also depend on the velocities of the particles, as well as on the interactions between particles. In section II we had already incorporated velocities, but in a classical context, where there was no stochastic component and the changes in particle numbers were due to being in contact with a reservoir. A natural question arises: what is the physical/mathematical connection between these approaches?
In this section, we focus on Langevin dynamics with varying copy numbers as a middle ground to hint at the similarities between the classical open systems approach from section II and the stochastic reaction-diffusion approach from section III.1. This will hint at a general approach to model the dynamics of systems with varying number of particles, enabling richer models of open systems.
@@ III.2 Langevin dynamics with varying number of particles
where simply denotes the infinitesimal generator of this Fokker-Planck equation for an particle system —in three dimensions, the equation would be a Fokker-Planck equation in dimensions. We would like to now incorporate processes that change the number of particles. This can be framed quite generally following section III.1 in terms of reactions. Reaction events can be thought of more generally as instantaneous events that follow a Poisson process in terms of rate functions . Note these rates must not be constant; they can depend on time, on the state of the system or on external variables. The master equation for this process will have the same form as the CDME in eq. 20, but now instead of the diffusion operator, we have the Klein-Kramers one
@@
In addition, the rate functions within the operators could also depend explicitly on the velocities of particles, allowing for richer particle-base models and generalizing the depth of field of the CDME. The approach presented here based on Langevin dynamics is an ideal starting ground to couple open particle-based models with fluids and to incorporate hydrodynamic and electrostatic interactions. As a final remark, note that if we take the overdamped limit of the Langevin dynamics, we in principle recover the CDME from section III.1.
IV A general master equation for systems with varying number of particles
In this section, we first show what are the mathematical connections between the approaches from sections II.1 and II.2 and the other two approaches in sections III.1 and III.2. Based on the insight gained in section III.2, we compare to the previous results and show how to formulate a general master equation for classical systems with varying number of particles.
In both previous sections II and III we introduced equations to describe the dynamics of densities in the phase space of systems with varying number of particles. The first insight is that mathematically the -particle Liouville equation is a special case of the particle Fokker-Planck equation (Klein-Kramers eq.). To show this, we start with the classical approach from section II.1. The resulting equation for the distribution dynamics in the phase space of the open system is given by a family of Liouville -particle equations coupled by the terms (eq. 10)
The Bergman-Lebowitz model from section II.2 shares the same mathematical structure but with different coupling terms due to different conceptual starting points of view: section II.1 employs an explicit particle reservoir whose coupling with the open system has been achieved by explicit integration of its degrees of freedom, while section II.2 employs a reservoir modeled a priori with a stochastic term for the exchange with the open system. In practice, the Bergman-Lebowitz model conceptually lies in between the other two approaches shown here. We can compare this equation with the th component of the master equation eq. 36 from section III.2
@@ IV A general master equation for systems with varying number of particles
The second insight is that these equations have an analogous mathematical structure. One part is essentially a transport term for the -particle density given by a Fokker-Planck equation, which simplifies to a Liouville equation in deterministic cases. The other part is a coupling term that models the change in the number of particles across the family of densities . We can write both equations in the following form:
@@
Section II.1, Liouville-like equation for an open subsystem:
@@
Section II.2, Bergman-Lebowitz equation of open systems:
(48) Once again the dynamics of the transport process are deterministic with Hamiltonian structure, so we can use again the Liouville operator. However, the interaction with the reservoir is modeled stochastically using the Markovian transition kernel . This can be understood under the same light as the reaction in section III.1. We can even separate the terms into a general loss and a gain term where
(49) (50) In this sense, this approach is a middle ground between the approach in section II.1 and the one in section III.1.
Section III.1, Chemical diffusion master equation for reaction diffusion:
@@
(52) Following [11], we can use the local rate functions for each reaction with these relations and write a global Markovian transition kernel in the form of eqs. 49 and 50, which establishes a direct connection with the Bergman-Lebowitz approach. The main difference is that in the Bergman-Lebowitz approach this kernel can depend on velocities while here it would only depend on positions.
Section III.2, Langevin dynamics with varying particle numbers:
@@
(53) From a mathematical point of view, this case covers the most general transport dynamics in this work; all the other cases are special cases of this one. Here, the transport dynamics are stochastic and track positions and velocities. Thus, the transport operator is given by the generator of the Klein-Kramers equation; the Fokker-Planck equation for Langevin dynamics. In the overdamped limit, we recover the diffusion operator from section III.1. Alternatively, if we remove the noise term, we recover the Liouville operator from sections II.1 and II.2. The coupling term works exactly as before, with the slight difference that now the reaction rate functions can also depend on the velocity. Following the same reasoning as before, we could once again write this in the form of eqs. 49 and 50, but unlike the overdamped Langevin case, we would also have a dependence on velocities.
@@
All the cases are special cases of eq. 46. However, its mathematical structure is even more general and does not need to be constrained to Langevin dynamics, Hamiltonian structure or a specific form of the coupling terms. The Liouville/Fokker-Planck term can be written for very general dynamical systems, and the coupling terms can model reactions, interactions with external reservoirs or any other kind of process that changes the number of particles. The only constraint is that every term in the equation must conserve the total probability. We know the Fokker-Planck equation is probability preserving, so we only need to make sure that also conserves the total probability (as formulated in eqs. 9, 19 and 37). This was shown in [12, 34] for the classical cases and in [8] for the reaction-diffusion case.
V Perspectives: physical and numerical

### v2

@@ III.2 The Bergman-Lebowitz equation of open systems
Bergmann and Lebowitz, and Lebowitz and Shimony in two seminal papers [51, 52], have proposed an equation based on the extension of the Liouville equation that models open systems embedded in a reservoir of particles and energy. The essence of the model is to consider an impulsive, Markovian interaction between the reservoir and the system; the reservoir is considered stationary and not influenced by the changes occurring in the system, thus the thermodynamic state point of the reservoir is fixed. The interaction between the system and the reservoir is modeled as a discontinuous transition of the system from a state with particles () to one with particles (), where corresponds to the canonical variables (position and momenta). The change of the state of the system state is described by a time-independent Markovian transition kernel, . The kernel, , corresponds to the probability per unit time that the system at makes a transition to as a result of the interaction with the reservoir. The probability density, , in some point of the phase space is then regulated by an extension of the Liouville equation:
@@
The first term of eq. 20 corresponds precisely to the Liouville equation for the particle system (eq. 11), while the second term establishes a coupling with the reservoir(s). This last term is composed of negative or positive contributions corresponding either to the outflux/loss of probability from the current state or to the influx/gain of probability into the current state, respectively.
@@
it follows that the stationary Grand Ensemble is the Grand Canonical ensemble; here where is the Boltzmann constant and the temperature, while is the chemical potential of the system. The key difference between the approach of section III.1 and the approach of Bergman and Lebowitz lies in the assumption about the reservoir and the corresponding term of system-reservoir exchange in the equation. The model of section III.1 does not require any assumption about the reservoir, but directly integrates its particle’s degrees of freedom. As a consequence, the system-reservoir coupling term is explicitly written in terms of particle quantities without any stochastic assumption. In the model shown here, the reservoir is modeled a priori without any explicit link to its particle resolution and the system-reservoir coupling term is modeled, as a consequence, with a probabilistic process. The model of Bergman and Lebowitz has been very important for the numerical implementation of the open system approach AdResS in an intermediate step. Such a model, differently from the model of section III.1, could not provide an explicit numerical receipt for the system-reservoir coupling, nevertheless it allowed for a physical qualitative interpretation of the coupling conditions of the AdResS code [2, 25].
IV Master equations for reaction-diffusion processes
@@
Molecular dynamics are limited to the study of complex biochemical phenomena at the scale of living cells. While the most sophisticated molecular simulations can achieve simulations of a few macromolecules in the scale of micro- or milliseconds when being optimistic, biochemical processes at these scales often involve thousands or millions of macromolecules and occur over timescales of seconds. Moreover, the chemical events at the molecular scale happen at a much faster scales than those relevant to life processes, and thus their detailed molecular kinetics do not play a key role in the dynamics. It is thus appropriate to consider particle-based reaction-diffusion models (fig. 2), where molecules are represented as particles undergoing random motion due to thermal fluctuations of the solvent (diffusion), and chemical reactions due to instantaneous reaction events, which often occur after chancy encounters between two or more molecules. The relevant features of the molecular scale are captured in the diffusion coefficients and the reaction rate functions.
@@
Particle-based reaction-diffusion models are often the standard model to describe biochemical processes at the scale of living cells. At these scales, the so-called chemical diffusion master equation (CDME) provides a probabilistic model in terms of the number of particles of each of the chemical species involved, as well as their spatial configuration [14, 18, 24]. One of the original objectives of developing the CDME was to serve as an underlying ground model upon which one can construct particle-based reaction-diffusion simulations consistently. Moreover, one can in principle recover most other reaction-diffusion models as limits of the CDME, e.g. in the well-mixed limit, one recovers the well-known chemical master equation (CME) [30, 34, 64, 68]; in the thermodynamic limit, one recovers reaction-diffusion PDEs; and when doing spatial discretizations, one recovers the reaction-diffusion master equation [41] or the spatiotemporal master equation [73]. Thus, it serves as a unifying framework from which one can develop numerical schemes that are consistent across scales. For instance, some mathematicians are interested in branching and annihilating Brownian motion [5] and its connection to the Kolmogorov-Petrovsky-Piskunov-Fisher (KPPF) equation. Following an applied mathematics/mathematical physics approach, one can frame these processes in terms of the CDME and study the limiting cases to recover the KPPF reaction-diffusion PDE, as done recently for other reaction systems [12]. This could yield alternative insights to the research community in that field, as well as methodologies for multiscale simulations.
@@
One of the main virtues of the CDME is its capability to handle systems with varying number of particles while maintaining spatial resolution. This is an inherent characteristic of the CDME since reaction events often change the number of particles in the system. In this section, we will overview the CDME for a simple example, and we will show how it extends when the diffusion term is replaced by Langevin dynamics.
a.
b.
@@
IV.1 Chemical diffusion master equation: coupling diffusion with reaction processes
@@
As a starting point, we follow [18] and consider a system with a varying number of particles, all corresponding to one chemical species, enclosed in a finite domain (with reflecting boundaries in the boundaries of ). The configuration of the system is given by the number of particles and their positions, so its probability distribution is given as an ordered family of probability density functions:
@@
The generator of the CDME is decomposed into the reaction and diffusion components: is the diffusion operator and corresponds to the reaction operator for the th reaction. Each reaction operator can be split into two contributions
representing the gain or loss of probability respectively due the th reaction in a given configuration.
To exemplify the formulation of the CDME for a reaction involving only one chemical species, we focus on the following reaction (fig. 2b.)
@@
In the absence of drift and assuming isotropic diffusion, would simply be the Laplacian, with a scalar diffusion constant . Considering this structure and assuming there are no reactions, the CDME would correspond to an infinite family of uncoupled Fokker–Planck equations for the particle positions, where each member of the family corresponds to a different number of particles.
@@
where the notation omitted the time dependence of for simplicity. The extension to multiple species is reported in the following subsection. The CDME for general one-species and multiple-species reactions were formulated in detail in [18, 14].
IV.1.1 Multiple species extension
@@
with rate function , where and are the locations of one pair of reactants and is the location of the product. The stochastic dynamics of the system are described in terms of the distributions , where indicate the numbers of , , and particles, respectively, and indicates the positions of the particles, of the particles, and of the particles. The normalization condition eq. 23 generalizes to multiple species
@@
where to simplify notation, the dependence of on time has been omitted, as well as its dependence on the positions if clear from context. The first line corresponds to the diffusion of each chemical species, the second line to the gain term and the last one to the loss. The works [18, 14] developed a comprehensive and formal mathematical framework to formulate the CDME for any given reaction system.
IV.2 Langevin dynamics with varying number of particles
There is a close mathematical link between section III and section IV.1; the resulting equations from both approaches describe the dynamics of densities in the phases space of systems with varying number of particles. In section IV.1, we introduced the diffusion operator, as well as the densities, only dependent on the position of the particles. However, one can in principle define these operators on more general dynamics. For instance, the state of the system can also depend on the velocities of the particles, as well as on the interactions between particles. In section III we had already incorporated velocities, but in a classical context, where there was no stochastic component and the changes in particle number were due to being in contact with a reservoir. A natural question arises: what is the physical/mathematical connection between these approaches?
In this section, we focus on Langevin dynamics with varying copy numbers as a middle ground to hint at the similarities between the classical open systems approach from section III and the stochastic reaction-diffusion approach from section IV.1. This will hint at a general approach to model the dynamics of systems with varying number of particles, enabling richer models of open systems.
@@ IV.2 Langevin dynamics with varying number of particles
where simply denotes the infinitesimal generator of this Fokker-Planck equation for an particle system with positions and velocities—in three dimensions, the equation would be a Fokker-Planck equation in dimensions. We would like to now incorporate processes that change the number of particles. This can be framed quite generally following section IV.1 in terms of reactions. Reaction events can be thought of more generally as instantaneous events that follow a Poisson process in terms of rate functions . Note these rates must not be constant; they can depend on time, on the state of the system or on external variables. The master equation for this process will have the same form as the CDME in eq. 24, but now instead of the diffusion operator, we have the Klein-Kramers one
@@
In addition, the rate functions within the operators could also depend explicitly on the velocities of particles, allowing for richer particle-base models and generalizing the depth of field of the CDME. The approach presented here based on Langevin dynamics is an ideal starting ground to couple open particle-based models with fluids and to incorporate hydrodynamic and electrostatic interactions. As a final remark, note that if we take the overdamped limit of the Langevin dynamics, we in principle recover the CDME from section IV.1.
V Previous models as special cases
In this section, we first show what are the mathematical connections between the approaches from sections III.1 and III.2 and the other two approaches in sections IV.1 and IV.2. Based on the insight gained in section IV.2, we compare to the previous results and show how this inspired the general master equation for classical systems with varying number of particles from section II.
In both previous sections III and IV, we introduced equations to describe the dynamics of densities in the phase space of systems with varying number of particles. The first insight is that mathematically the -particle Liouville equation is a special case of the particle Fokker-Planck equation (Klein-Kramers eq.). To show this, we start with the classical approach from section III.1. The resulting equation for the distribution dynamics in the phase space of the open system is given by a family of Liouville -particle equations coupled by the terms (eq. 16)
The Bergman-Lebowitz model from section III.2 shares the same mathematical structure but with different coupling terms due to different conceptual starting points of view: section III.1 employs an explicit particle reservoir whose coupling with the open system has been achieved by explicit integration of its degrees of freedom, while section III.2 employs a reservoir modeled a priori with a stochastic term for the exchange with the open system. In practice, the Bergman-Lebowitz model conceptually lies in between the other two approaches shown here. We can compare this equation with the th component of the master equation eq. 40 from section IV.2
@@ V Previous models as special cases
The second insight is that these equations have an analogous mathematical structure. One part is essentially a transport term for the -particle density given by a Fokker-Planck equation, which simplifies to a Liouville equation in deterministic cases. The other part is a coupling term that models the change in the number of particles across the family of densities . We can write both equations in the form of eq. 5:
@@
Section III.1, Liouville-like equation for an open subsystem:
@@
Section III.2, Bergman-Lebowitz equation of open systems:
(48) Once again the dynamics of the transport process are deterministic with Hamiltonian structure, so we can use again the Liouville operator. However, the interaction with the reservoir is modeled stochastically using the Markovian transition kernel . This can be understood under the same light as the reaction in section IV.1. We can even separate the terms into a general loss and a gain term where
(49) (50) In this sense, this approach is a middle ground between the approach in section III.1 and the one in section IV.1.
Section IV.1, Chemical diffusion master equation for reaction diffusion:
@@
(52) Following [18], we can use the local rate functions for each reaction with these relations and write a global Markovian transition kernel in the form of eqs. 49 and 50, which establishes a direct connection with the Bergman-Lebowitz approach. The main difference is that, in the Bergman-Lebowitz approach, this kernel can depend on velocities while here it would only depend on positions.
Section IV.2, Langevin dynamics with varying particle number:
@@
(53) From a mathematical point of view, this case covers the most general transport dynamics from all the examples presented; all the other cases are special cases of this one. Here, the transport dynamics are stochastic and track positions and velocities. Thus, the transport operator is given by the generator of the Klein-Kramers equation; the Fokker-Planck equation for Langevin dynamics. In the overdamped limit, we recover the diffusion operator from section IV.1. Alternatively, if we remove the noise term, we recover the Liouville operator from sections III.1 and III.2. The coupling term works exactly as before, with the slight difference that now the reaction rate functions can also depend on the velocity. Following the same reasoning as before, we could once again write this in the form of eqs. 49 and 50, but unlike the overdamped Langevin case, we would also have a dependence on velocities.
@@
All the cases are special cases of eq. 46. However, its mathematical structure is even more general and does not need to be constrained to Langevin dynamics, Hamiltonian structure or a specific form of the coupling terms. The Liouville/Fokker-Planck term can be written for very general deterministic or stochastic dynamical systems, and the coupling terms can model reactions, interactions with external reservoirs or any other process that changes the number of particles. The only constraint is that every term in the equation must conserve the total probability. We know the Fokker-Planck equation part is probability preserving, so we only need to make sure that the couplings also preserve the total probability ( eqs. 15, 23 and 41) following eq. 6. This was shown in [20, 51] for classical molecular dynamics and in [14] for reaction-diffusion dynamics. For other specific applications, one must construct the corresponding transport and coupling operators.
VI Perspectives: physical, numerical and beyond
The general formulation of the master equation for systems with varying particle number presented in this work provides a novel and synergistic link between different fields and how they handle open settings. Through its general formulation, it unveils new perspectives and opens up the application scope to a diverse range of fields within and beyond physics. There is ample potential for future applications, specifically in the design of multiscale theory and simulation of complex systems. In this section, we discuss new perspectives inspired by this work, as well as current and future applications.